# Plasmonic metamaterial time crystal

### v1

@@ Methods
where $\mathbf{G}(\tau)$ and $\mathbf{\mu}$ are both $2\times 2$ matrices acting on the spinor space of the cavity mode (a^,a^)($\hat$$^{\dagger},\hat{a}$) (see supplementary material). $\mathbf{G}(\tau)$ is periodic with the same period as the drive and describes the periodic micro-motion of the system while the Floquet exponent matrix $e^{-\mathrm{i}\mathbf{\mu}t}$ describes its long-term behavior. In the absence of dissipation, the imaginary part of the eigenvalues $\mu_{1,2}$ of matrix $\mathbf{\mu}$ determine wether the system is stable ($\mathrm{Im}(\mu_{1,2})<0$) or unstable ($\mathrm{Im}(\mu_{1,2})>0$), as quantities evolve as $e^{\mathrm{Im}(\mu_{1,2})t}$ at long times. In the presence of dissipation characterized by a total decay rate $\Gamma=\frac{\gamma+\gamma^{\prime}}{2}$, it can be shown that eigenvalues of $\mathbf{\mu}$ undergo a shift by $\Gamma$ while $\mathbf{G}(t)$ remains unaffected. Defining $\mathrm{Im}(\mu)=\max\limits_{1,2}\mathrm{Im}(\mu_{1,2})$, this allows to conclude directly on the stability of the dynamics from a contour line defined by $\frac{|\mathrm{Im}(\mu)|}{\omega_{\mathrm{c}}^{0}}=\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$ and which marks the transition between stable ($\frac{|\mathrm{Im}(\mu)|}{\omega_{\mathrm{c}}^{0}}<\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$) and unstable dynamics ($\frac{|\mathrm{Im}(\mu)|}{\omega_{\mathrm{c}}^{0}}>\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$). The total normalized decay rate $\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}=0.12$ presented in Fig. 3a. corresponds to the experimentally measured one and can be found in Table 1.
Wigner function (Figure 3b)
The Wigner function distribution of the cavity state, plotted in Fig. 3b., is defined via its density matrix $\hat{\rho}(\tau)$ as :
$W(\alpha,\tau)=\frac{2}{\pi}\text{Tr}\left(\hat{\rho}(\tau)\hat{D}(\alpha)(-1)^{a^{\dagger}a}\hat{D}(-\alpha)\right),$ (12)
where for all $\alpha\in\mathbb{C}$, the displacement operator is defined by $\hat{D}(\alpha)=\exp(\alpha a^{\dagger}-\alpha^{\ast}a)$. It is computed from the time evolution of the density matrix according to its Hamiltonian eq.(7) and accounting for losses at the rate $\Gamma=\frac{\gamma+\gamma^{\prime}}{2}$ as :
$\partial_{t}\hat{\rho}=-\mathrm{i}\left[\hat{H}(\tau),\hat{\rho}\right]+\frac{\gamma+\gamma^{\prime}}{2}\mathcal{D}[a]\hat{\rho}$. (13)
where $\mathcal{D}[a]$ is the the one-photon loss dissipator which acts on the density matrix through $\mathcal{D}[a]\rho=a\rho a^{\dagger}-\frac{1}{2}\left\{a^{\dagger}a,\rho\right\}$. Further details of the calculations are given in the supplementary materials.
Photonic time crystal (Figure 3c)
To treat the case of photonic time crystals, we extend the formulation of the model to general dispersive systems. Eigenmodes depend on index $k$, which corresponds either to the parallel wavevector (translationally-invariant systems) or Bloch-wavevector (periodically structured metamaterials) of a given eigenmode. For such systems, the Hamiltonian writes (see supplementary materials):
$\hat{H(\tau)}=\sum_{k}\hbar(\omega_{\rm c}^{0}(k)+\eta_{k}(\tau))\hat{a}_{k}^{\dagger}\hat{a}_{k}-\frac{\hbar\eta_{k}(\tau)}{2}\left(\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}+{\hat{a}}_{k}{\hat{a}}_{-k}\right)$ (14)
where $\omega_{\rm c}^{0}(k)$ and $\eta_{k}(\tau)$ are, respectively, the $k$-dependent equilibrium resonance frequency and modulation parameter of the eigenmodes. As we consider systems that are driven homogeneously at $k=0$, creation and annihilation of plasmon pairs have opposite momenta in order to satisfy momentum conservation rules.
Computation of $\omega_{\rm c}^{0}(k)$ and $\eta_{k}(\tau)$ is performed with an identical procedure as described above for a single cavity: the dispersion relation of the system is computed both at equilibrium and through variation of the effective mass that enters the plasma frequency, allowing for the determination of the relation of proportionality as in eq. 9 between $\frac{\eta_{k}(\tau)}{\omega_{\rm c}^{0}(k)}$ and $\frac{\delta m^{*}(\tau)}{m_{0}^{*}}$ for any given $k$-mode. Floquet analysis is then performed on Hamiltonian eq. 14 in order to determine the Floquet eigenvalues $\omega_{1,2}$ of the driven system and to track the opening of $k$-gaps. After accounting for non-radiative losses with a shift in the imaginary part and normalizing by the equilibrium plasma frequency $\omega_{\mathrm{p}}$, real and imaginary parts are reported in Fig. 3c in the two following cases for: 1) a dispersive insulator/plasmonic interface and 2) our dispersion-free metamaterial design for cavities tuned in the electrostatic regime.

### v2

@@ Methods
In the absence of dissipation, the imaginary part of the eigenvalues $\mu_{\pm}$ determine wether the system is stable ($\mathrm{Im}(\mu_{\pm})<0$) or unstable ($\mathrm{Im}(\mu_{\pm})>0$), as quantities evolve as $e^{\mathrm{Im}(\mu_{\pm})\tau}$ at long times. In the presence of dissipation characterized by a total decay rate $\Gamma=\frac{\gamma+\gamma^{\prime}}{2}$, it can be shown that eigenvalues of $\mathbf{\mu}$ undergo a shift by $i\Gamma$ while $\mathbf{G}(t)$ remains unaffected. This allows to conclude directly on the stability of the dynamics from a contour line defined by $\frac{|\Delta\Gamma|}{\omega_{\mathrm{c}}^{0}}=\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$ and which marks the transition between stable ($\frac{|\Delta\Gamma|}{\omega_{\mathrm{c}}^{0}}<\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$) and unstable dynamics ($\frac{|\Delta\Gamma|}{\omega_{\mathrm{c}}^{0}}>\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}$). The total normalized decay rate $\frac{\Gamma}{\omega_{\mathrm{c}}^{0}}=0.12$ presented in Fig. 4b. corresponds to the experimentally measured one and can be found in Table 1.